### v1

# thanks: Contributed equallythanks: Contributed equally

Abstract
Continuous-time quantum walks (CTQWs) on dynamic graphs, referred to as dynamic CTQWs, are a recently introduced universal model of computation that offers a new paradigm in which to envision quantum algorithms. In this work we develop a mapping from dynamic CTQWs to the gate model of computation in the form of an algorithm to convert arbitrary single edge walks and single self loop walks, which are the fundamental building blocks of dynamic CTQWs, to their circuit model counterparts. We use this mapping to introduce an arbitrary quantum state preparation framework based on dynamic CTQWs. Our approach utilizes global information about the target state, relates state preparation to finding the optimal path in a graph, and leads to optimizations in the reduction of controls that are not as obvious in other approaches. Interestingly, classical optimization problems such as the minimal hitting set, minimum spanning tree, and shortest Hamiltonian path problems arise in our framework. We test our methods against uniformly controlled rotations methods, used by Qiskit, and find ours requires fewer CX gates when the target state has a polynomial number of non-zero amplitudes.
preprint: APS/123-QED
I Introduction
Quantum algorithms have the potential to provide speed-up over classical algorithms for some problems [1, 2]. However, certain quantum algorithms may require non-trivial input states [3, 4], which in general, are challenging to prepare and may require exponentially many CX gates [5] in the worst case. Nevertheless, many practically relevant quantum states can be prepared much more efficiently by taking advantage of their special properties or their sparsity [6]. In this work we focus on the latter, and show how continuous-time quantum walks (CTQWs) on dynamic graphs can be used to create arbitrary sparse (and dense) quantum states.
CTQWs on graphs are a universal model of computation [7] that excels at spatial searches [8, 9] and has applications in finance [10], coherent transport on networks [11], modeling transport in geological formations [12], and combinatorial optimization [13]. In this model, the quantum state vector is evolved under the action of $e^{-iAt}$ for some time $t$, where $A$ is an adjacency matrix of some fixed unweighted graph. Such walks have been implemented natively on photonic chips [14, 15, 16].
In 2019, CTQWs on dynamic graphs (i.e. graphs that may change as a function of time) were introduced and shown to also be universal for computation [17] by implementing the universal gate set H, CX, and T.
In the original dynamic graph model, isolated vertices were propagated as singletons. However a new model where isolated vertices are not propagated was introduced later in Ref. [18]. Simplification techniques were introduced that can reduce the length of the dynamic graph sequence if graphs in the sequence satisfy certain properties [19]. Furthermore, the authors of [20] found that CTQWs on at most three dynamic graphs can be used to implement the equivalent of a universal gate set.
However, the opposite conversion from CTQWs to the gate model is less studied. Since CTQWs on dynamic graphs have yet to be implemented on hardware, converting them to the gate model is necessary to simulate them on existing hardware. Additionally, developing an algorithm for such a conversion offers an alternative way of thinking about circuit design, which may result in simpler circuits, compiler optimization techniques, and new ansatzes for QAOA [21, 22, 23, 24], MA-QAOA [25, 26, 27] or VQE [28, 29, 30], where the trainable parameters correspond to dynamic graph propagation times.
To this end, we develop an algorithm that converts CTQWs on single edge graphs and single self-loop graphs, which can serve as a basis for arbitrary graphs, to a sequence of gates in the circuit model. This conversion algorithm is used as a foundation for a new deterministic arbitrary quantum state preparation (QSP) method, where the CX count of the resulting circuit is linear in the number of $m$ non-zero amplitude computational basis states that comprise the target state and the number of qubits $n$. Fig. 1 gives a general picture of the quantum walks QSP framework introduced in this paper. Our framework does not require ancillas.
Figure 1: This figure lays out the framework for QSP via quantum walks. Alternating single edge quantum walks with self-loop quantum walks, allows arbitrary QSP. We provide conversion methods to construct the gate based circuit. Control reduction is a method for reducing controls that we introduce in this paper.
Our method builds on the idea of exploiting sparsity to increase the efficiency of QSP techniques by approaching the problem from the perspective of dynamic quantum walks. As such, our method is best suited for the asymptotically sparse states ($m=\mathcal{O}(poly(n))$), but can, in principle, also work for the asymptotically dense states ($m=\mathcal{O}(2^{n})$). The walk framework that we present here is flexible, scalable, and intuitive, building upon well-established graph-based algorithms to efficiently prepare arbitrary quantum states.
This paper is organized as follows. First, we introduce necessary background information and previous works in QSP in Sec. II. We then introduce an algorithm that converts CTQWs on single edge dynamic graphs and single self loop graphs to the circuit model in Sec. III. In Sec. IV, we introduce a deterministic state preparation algorithm and analyze its circuit complexity. We compare our state preparation methods to the uniformly controlled rotation method [5, 31] used by Qiskit in Sec. V. Finally, we discuss future work in Sec. VI.
II Background
In this section, we provide background information and examples of CTQWs on dynamic graphs and state-of-the-art state preparation circuits.
II.1 Continuous-Time Quantum Walks on Dynamic Graphs
A dynamic graph is defined as a sequence of ordered pairs $\{(G_{i},t_{i})\}_{i=1}^{\ell}$ for some $\ell\in\mathbb{N}$ that consists of unweighted graphs $G_{i}$ and corresponding propagation times $t_{i}$. A CTQW on a dynamic graph is then a CTQW where the first walk is performed on graph $G_{1}$ for time $t_{1}$, followed by a walk on graph $G_{2}$ for time $t_{2}$, and so on until all graphs in the sequence have been walked upon. The final state |$\ket{\psi_{\ell}}$ is then given by
|$\ket{\psi_{\ell}}=e^{-iA_{\ell}t_{\ell}}\ldots e^{-iA_{2}t_{2}}e^{-iA_{1}t_{1}}\ket{\psi_{0}}.$ (1)
where $A_{i}$ is the adjacency matrix for graph $G_{i}$, that is a 0-1 valued symmetric matrix. For simplicity, we assume that each graph has $2^{n}$ vertices for some $n\in\mathbb{N}$, and each vertex represents a computational basis state.
As an example, consider the dynamic graph found in Fig. 2, where the adjacency matrices are
$A_{1}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}$
$A_{2}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}$
Figure 2: Dynamic CTQW implementation of CX gate.
If the walker starts in the initial state |$\ket{\psi_{0}}=(a,b,c,d)^{T}$, then the final state of this walker is
|$\ket{\psi_{2}}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&i&0\\ 0&0&0&i\end{pmatrix}\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&-i\\ 0&0&-i&0\end{pmatrix}\begin{pmatrix}a\\ b\\ c\\ d\end{pmatrix}=\begin{pmatrix}a\\ b\\ d\\ c\end{pmatrix}$
which is equivalent to a CX gate in the circuit model.
Interestingly, CTQWs on dynamic graphs that correspond to operations in the gate model tend to have a similar form: some phase is added to particular basis states via self-loops on vertices of the graph, then they are connected by edges to allow for state transfer, followed by additional self-loops to eliminate unwanted phase. This general alternating sequence of graphs inspires the state preparation algorithm introduced later in this work.
II.2 Prior State Preparation Results
Due to its fundamental role in quantum algorithms [32, 33, 34], quantum state preparation is an active area of research within the space of quantum computation. The resource overhead of general QSP problems for arbitrary states is known to be $\mathcal{O}(2^{n})$ in the number of CX gates [5, 31]. Within this limit, there are different techniques utilizing different resources, which can be useful depending on the context. For instance, if depth is more of a concern than space, Ref. [35] provides an $\mathcal{O}(n^{2})$ depth divide-and-conquer method to prepare arbitrary states at the cost of needing $\mathcal{O}(2^{n})$ ancillary qubits. For arbitrary QSP, the state of the art deterministic protocol achieves $\mathcal{O}(\frac{23}{24}2^{n})$ CX scaling on even numbers of qubits [36]. The downside of this method is that it relies on the Schmidt decomposition of the $n$-qubit target state, a computationally expensive task, which has a classical run time of $\mathcal{O}(2^{3n/2})$ [37].
While arbitrary state preparation is exponential, there are states of practical interest that do not require exponential resources, even in the worst case [38, 39]. Manual methods are strategies for QSP that use advanced knowledge of the target state to save resources in ways that might not be obvious when preparing arbitrary states. For example, GHZ states are highly entangled, but only require a linear number of CX gates. Similarly, the methods introduced in [40] require $\mathcal{O}(kn)$ CX gates to prepare $n$-qubit Dicke states of Hamming weight $k$. For machine learning applications on classical data, quantum data loaders can prepare sparse amplitude encodings of $d$-dimensional real-valued vectors on $d$-qubits in $\mathcal{O}(\log d)$ circuit depth [41]. In Ref. [42], Zhang et al. introduce a technique to prepare input states to the HHL algorithm [3] based on the finite element method equations relevant to electromagnetic problems. Zhang et al.’s algorithm achieves $\mathcal{O}(n)$ depth scaling by exploiting the symmetry and sparsity of the relevant target states; however, these useful assumptions are not guaranteed for arbitrary states.
The state of the art method for preparing arbitrary sparse states is detailed in [6], which produces circuits with $\mathcal{O}(nm)$ CX gates and runs in $\mathcal{O}(nm^{2}\log(m))$ time classically. This type of method is attractive for real-world applications of quantum computers because they can take advantage of sparsity while still remaining agnostic to particular details of the state. Another method using decision trees [43] has $\mathcal{O}(nm)$ CX gates, but it requires an ancilla qubit. The method in Ref. [44] also achieves $\mathcal{O}(nm)$ CX gates. However, this method is based on Householder reflections and is more complicated than the one introduced in Ref. [6].
III Conversion from CTQW on dynamic graphs to the Circuit Model
A standard basis for any adjacency matrix is given by the set of adjacency matrices corresponding to all single edge walks and all single self-loop walks. We introduce the conversion methods that can construct the gate-based representations for these walks on $n$ qubits. Throughout this paper, the considered states are represented in the standard computational basis set.
III.1 Single Edge CTQW to Circuit Model
The adjacency matrix for a single edge graph connecting basis states |$\ket{j}$ and |$\ket{k}$ is given by
$A(j,k)=\ket{j}\!\!\bra{k}+\ket{k}\!\!\bra{j}$ | (2)
and the corresponding CTQW for time $t$ is
[…]
This unitary transfers amplitude between states |$\ket{j}$ and |$\ket{k}$ and leaves all other states untouched.
If |$\ket{j}$ and |$\ket{k}$ differ in exactly one bit at position $l$, $U(j,k;t)$ is the $(n-1)$-controlled Rx($2t$) gate
$\text{Rx}(2t)=\begin{pmatrix}\cos(t)&-i\sin(t)\\ -i\sin(t)&\cos(t)\end{pmatrix}$ (4)
where the target qubit is $l$, and the remaining qubits are either 0- or 1-controls corresponding to the remaining bits of |$\ket{j}$ and |$\ket{k}$.
If the Hamming distance between |$\ket{j}$ and |$\ket{k}$ is greater than $1$, using the CRx($2t$) gate is still possible, but requires some extra overhead. First apply a sequence of CX gates where the control is any bit in which |$\ket{j}$ and |$\ket{k}$ differ and the targets are the other bits in which |$\ket{j}$ and |$\ket{k}$ differ. Then apply the CRx($2t$) gate as before and apply the reverse of the conjugating CX sequence. An example conversion of a single edge walk to gates is shown in Fig. 3. The first CX gate transforms the two basis states such that their Hamming distance becomes equal to 1. The CRx creates the superposition and the final CX restores the original basis states.
The $(n-1)$-controlled CRx gate is a multi-controlled special unitary, and can be decomposed in $\mathcal{O}(n)$ CX and single-qubit gates [45].
Figure 3: When a single edge walk is between states with non-unit Hamming distance from each other, CX gates might be required in addition to the CRx gate.
III.2 Single Self-Loop CTQW to Circuit Model
A single self-loop walk on basis state |$\ket{j}$ is given by the adjacency matrix
$A(j)=\outerproduct{j}{j},$ (5)
The corresponding CTQW for time $t$ is
$U(j;t)=e^{-itA(j)}=e^{-it}\outerproduct{j}{j}+\sum_{k\neq j}\outerproduct{k}{k},$ (6)
which is a diagonal matrix with a single non-unit element on the diagonal.
From this equation, one can see that the self-loop walk on a single vertex is equivalent to adding phase to exactly one computational basis state. This operation corresponds to an application of an $(n-1)$-controlled phase gate P for time $-t$
$\text{P}(t)=\begin{pmatrix}1&0\\ 0&e^{it}\end{pmatrix},$ (7)
where the target qubit can be arbitrarily chosen among the bits of $j$ that are equal to 1 (or 0, with the appropriate X conjugation), and the remaining qubits are either 0- or 1-controls corresponding to the remaining bits of |$\ket{j}$.
However, the $\text{P}(t)$ gate is not a special unitary. As such, without using ancilla qubits, its decomposition requires $\mathcal{O}(n^{2})$ CX gates [46]. A more efficient (but also more nuanced) approach is to use an $(n-1)$-controlled Rz($2t$) gate
$\text{Rz}(2t)=\begin{pmatrix}e^{-it}&0\\ 0&e^{it}\end{pmatrix}$ (8)
which is a special unitary and can be decomposed in $\mathcal{O}(n)$ CX gates [45].
At a first glance, an $(n-1)$-controlled Rz gate affects two adjacent computational basis states, which makes it not equivalent to a self-loop walk in general. However, if only one of these two basis states exists in the state affected by the CRz gate, then it becomes essentially equivalent to the CP gate and can also be used to implement a self-loop walk.
As long as the state we apply the walk to (|$\ket{\psi_{0}}$ in Eq. 1) has at least one zero-amplitude basis state |$\ket{z}$, not necessarily adjacent to |$\ket{j}$, one can use the same CX conjugation technique as described in the previous section for the Rx gate. This enables the interaction between |$\ket{j}$ and |$\ket{z}$ via the CRz gate and implements the self-loop walk on |$\ket{j}$.
In the case when |$\ket{\psi_{0}}$ is fully dense, i.e. all $2^{n}$ basis states have non-zero amplitudes, the above method will not work and a CP gate will have to be used instead.
III.3 Universal Computation
As it was mentioned in the Introduction, CTQWs on dynamic graphs are known to be universal [17]. However, the authors of Ref. [17] use walks on arbitrary graphs in their proof of universality. In this section we prove that using only single edge and single self-loop walks is sufficient to decompose an arbitrary unitary.
Proposition 1.
An arbitrary $d\times d$ unitary can be decomposed into a series of single self-loop and single edge CTQWs.
Proof.
The case of $d=1$ is trivial so we consider $d\geq 2$. An arbitrary unitary can be decomposed into a series of 2-level unitaries (unitaries that act non-trivially on two or fewer basis states, pages 189-191 of [47]).
Similarly to how an arbitrary single-qubit unitary can be decomposed as $U=W(\alpha)\text{Rz}(\beta)\text{Rx}(\gamma)\text{Rz}(\delta)$ (page 175 of [47]), where $W=e^{i\alpha}I$, an arbitrary 2-level unitary $U$ can be decomposed as
$U=\text{W'}(\alpha)\text{Rz'}(\beta)\text{Rx'}(\gamma)\text{Rz'}(\delta)$ (9)
for some real numbers $\alpha,\beta,\gamma,\delta$, where the W’, Rz’ and Rx’ are 2-level unitaries whose action in the corresponding 2D subspace is equivalent to their single-qubit counterparts.
As it was shown in Sections III.1 Eq. (3) and III.2 Eq. (6), a single edge CTQW corresponds to Rx’, and a sequence of two single self-loop CTQWs corresponds to W’ and Rz’. Thus, an arbitrary unitary can be decomposed into a series of single self-loop CTQWs and single edge CTQWs.
IV Deterministic state preparation via quantum walks
The above conversions allow one to write single edge walk dynamics and single self-loops in terms of the circuit model. In this section, we describe how a series of such quantum walks can be used to prepare an arbitrary quantum state.
IV.1 General Approach
Given the task of preparing a quantum state with $m<2^{n}$ non-zero amplitudes on $n$ qubits, we start by presenting a high-level description of the algorithm to accomplish this task:
  1. 1.
    Create a tree graph connecting the non-zero amplitudes, as shown in Fig. 4.
  2. 2.
    Starting from an arbitrarily chosen root, and following a graph traversal order, perform a self-loop walk for each encountered node and a single edge quantum walk for each encountered edge.
  3. 3.
    Convert the sequence of the quantum walks to the circuit representation.
Initially, the system starts with all amplitude in the root state, which can be constructed from the ground |$\ket{0}$ state simply by applying the X gates to the necessary qubits. The graph traversal order can be be determined by any graph traversal algorithm, such as Depth First Search or Breadth First Search.
An arbitrary amount of amplitude (absolute value) from the source can be transferred to the destination for any pair of the connected states corresponding to each single edge walk, and all non-zero amplitude states are connected transitively via a tree. Therefore, an arbitrary distribution of absolute values of amplitude (but not phases) can be established on the set of non-zero amplitude basis states via single edge walks.
Each self-loop walk on a basis state |$\ket{z_{i}}$ establishes the correct phase of the corresponding coefficient $c_{i}$, but does not change the absolute value of it. Therefore, performing self-loop walks on each non-zero amplitude basis state establishes the correct phases of the corresponding coefficients, without interfering with the absolute values established by the single edge walk. Therefore, an arbitrary quantum state can be prepared with a sequence of single edge walks and self-loop walks, as described in the algorithm above.
In the context of the state preparation, CRz gates can easily be used instead of CP gates to implement the self-loop walks, since every single edge walk transfers the amplitude to a new basis state with zero amplitude. Thus, as described in the previous section, the same basis state can be used to establish the correct phase on the source state for the single edge walk without additional CX conjugations. However, this will not apply to the leaf nodes (degree one vertices) in the tree, which will require another zero-amplitude state to interact with.
The tree contains $\mathcal{O}(m)$ nodes and edges, therefore $\mathcal{O}(m)$ walks will be required. As described in the previous section, each walk is represented by a single multi-controlled gate (CRz or CRx) that has up to $n-1$ controls. Each such gate can be implemented in $\mathcal{O}(n)$ CX gates [45] and may need to be conjugated by $\mathcal{O}(n)$ additional CX gates due to the Hamming distance between the adjacent basis states. Therefore the overall complexity of the algorithm, in terms of the required number of CX gates, is $\mathcal{O}(nm)$.
Figure 4: Although there are exponentially many basis states in a generic quantum state, when only a few have non-zero amplitude, a tree over the relevant states can be found with reasonable overhead.
IV.2 Control Reduction
The number of CX gates necessary to implement the multi-controlled Rz and Rx gates is proportional to the number of controls on those gates. As mentioned earlier, we might need up to $n-1$ controls in the worst case, but depending on the walk that we want to implement and the basis states that we have already visited, we might need less than that. For example, when we perform the first walk from the root, we never need any controls, since there are no other basis states that would be affected by the Rx gate.
More generally,
Proposition 2.
Let $n$ denote the number of qubits and $S=\{\ket{z_{1}},\ket{z_{2}},...,\ket{z_{k}}\}$ denote the set of visited nodes. Suppose we wish to perform a single edge walk from |$\ket{z_{j}}\in S$ to |$\ket{z_{\ell}}\notin S$ and |$\ket{z_{j}}$ and |$\ket{z_{\ell}}$ have a Hamming distance of 1. Let $b$ denote the qubit where |$\ket{z_{j}}$ is different from |$\ket{z_{\ell}}$, i.e. $z_{j}[b]\neq z_{\ell}[b].$ Let $D=\{\{k\,|\,z_{i}[k]\neq z_{j}[k],k\neq b\}\,|\,z_{i}\in S\setminus z_{j}\}$ denote the set of differing bits of the visited nodes with $z_{j}$ excluding the qubit $b$. For the gate-based representation of a single edge walk from basis state |$\ket{z_{j}}$ to |$\ket{z_{\ell}}$, it is sufficient to control the CRz or CRx gate on any hitting set of $D$, where the values of controls are equal to the corresponding bits of $z_{j}$. A hitting set of $D$ is a collection of elements $h$ such that $h\cap d_{i}\neq\emptyset$ for all $d_{i}\in D$.
Proof.
Adding a control on an arbitrary qubit $c$ to any gate makes the gate act only on the states conforming to the value of that control, i.e. |$\ket{z_{k}}$ such that $z_{k}[c]=0$ or 1, depending on the state of the control. By the definition of $d_{i}\in D$, controlling the CRz or CRx gate on any index $e\in d_{i}$ with the value of control equal to $z_{j}[e]$ will ensure that the gate does not act on $z_{i}$. Thus, controlling on any hitting set of $D$ will ensure that that the gate does not act on all $z_{i}\in S$, except $z_{j}$. ∎
When $z_{j}$ and $z_{\ell}$ have a Hamming distance greater than 1, we first have to update $S\rightarrow\tilde{S}$ with the conjugating CX gates, as described in Sec. III.1. Then we apply the previously described control reduction method to $\tilde{S}$.
In order to minimize the number of controls, one needs to minimize the size of the hitting set, i.e. solve the minimum hitting set problem, which is known to be NP-complete. As such, no polynomial algorithm is known to solve it exactly, but a number of heuristics exist that can provide good suboptimal solutions quickly [48, 49]. In practice, solving this problem for each walk drastically reduces the number of control qubits on many gates in the sequence (see Fig. 5).
Figure 5: The average number of CX gates necessary to implement a random state with $n$ qubits with and without control reduction. Each state consist of $m=n$ non-zero amplitudes and each point is the average over 1000 random states.
IV.3 Walk Order
Any tree constructed in the first step is, in principle, sufficient for state preparation but not all trees result in the same CX gate count. As mentioned in the previous section, the cost of a walk between a given pair of basis states depends on the Hamming distance between them. Therefore, connecting the basis states that are close to one another (in terms of their Hamming distance) minimizes the number of CX gates in the resulting circuit. Let us consider some of the more advantageous options.
Minimum Spanning Tree. The most straightforward option to minimize the Hamming distance between the connected basis states is to build a Minimum Spanning Tree (MST) over the complete graph of target basis states, where each edge is weighted by the Hamming distance between its endpoints.
The cost of calculating all pairwise Hamming distances is $\mathcal{O}(nm^{2})$, while the MST itself can be built in $\mathcal{O}(m^{2})$ with Prim’s algorithm. Thus, the overall classical complexity for this method is $\mathcal{O}(nm^{2})$.
In the case when the target state is fully dense, i.e. $m=2^{n}$, this method can be simplified, since the tree can be automatically built without calculating all pairwise distances by branching off sequentially in each dimension of the hypercube (see Fig. 6b). The same method can also be used for generally dense states, i.e. when $m=\mathcal{O}(2^{n})$. In this case, the resulting tree will go through some zero-amplitude states, which is not optimal for the circuit, but allows to save classical computational resources.
Shortest Hamiltonian Path. MST approach minimizes the total Hamming distance for the walks, which corresponds to the optimal walk order if control reduction is not taken into account. However, when control reduction is applied, other walk orders may be more efficient since they may allow for additional control reduction.
In general, it is difficult to calculate in advance which walk orders are better for the maximum control reduction, but from the numerical experiments we discovered that, on average, path graphs (trees with exactly two leaves) are more amenable to it than other trees. Therefore, instead of MST, one could find the Shortest Hamiltonian Path (SHP) in the same graph of the target basis states as used to build the MST.
This option is more computationally expensive, since the cost of finding SHP exactly is $\mathcal{O}(m!)$. However, approximate SHP heuristics can find suboptimal solutions much faster [50].
Similar to MST, when $m=2^{n}$ the procedure can be simplified by building a Hamiltonian path throughout the whole hypercube of basis states without calculating all pairwise distances. This path is the known as the Gray code. Examples of the quantum walks produced by SHP for the cases of sparse and dense states are shown in Fig. 6.
Sorted order. Another method to build a simple linear path without solving SHP is to connect the basis states sequentially in increasing order. This method is less optimal than SHP, but it is computationally inexpensive since the target basis states can be sorted in $\mathcal{O}(nm\log(m))$, and can provide us with a reasonably good walking order. For example, for a fully dense state, the average Hamming distance given by the sorted walk order is 2, which is much better than $n/2$ achieved by a random walk order.
A comparison of the different walking orders presented in this section for the case of $m=n$ with control reduction is shown in Fig. 7. We empirically validate the four quantum walk order methods and find that on the prepared states, SHP requires the fewest CX gates, although it approaches the MST method as $n$ increases.
(a) MST for a sparse state.
(b) MST for a dense state.
(c) SHP for a sparse state.
(d) SHP for a dense state.
Figure 6: Example walks generated by different methods for the case of 3 qubits. The weight of an edge is the Hamming distance of the two nodes. A potential walk path chosen by a given method is shown in red. In Figs. 6d and  6b, the costs of the edges are suppressed except for the path.
Figure 7: The average number of CX gates necessary to implement a random state with given number of qubits for the considered walk orders. Each state consist of $m=n$ non-zero amplitudes and each point is the average over 1000 random states.
IV.4 Additional Optimizations
A single edge between states |$\ket{j}$ and |$\ket{k}$ corresponds to a CRx gate, which introduces an imaginary phase on |$\ket{k}$. However, for the purpose of preparing a state with some real-valued amplitudes, a better approach is to use a CRy gate instead, which has the same action as CRx, except it does not introduce a complex phase. CRy keeps the amplitudes real and makes it unnecessary to use CRz or CP gates when the amplitude is transferred between the basis states with real coefficients. As such, for these states, one might want to use a special walk order that goes through all real-valued basis states before moving to the complex-valued ones. For the states where all amplitudes are real, this approach can reduce the total number of CX gates in the circuit by a factor of approximately 2. We plan to implement this in our code in the future.
IV.5 Example
Suppose we want to prepare the state
[…]
Using the algorithm described above, with sorted walk order and control reduction, one can generate a circuit shown in Fig. 8. The exact values of the coefficients $c_{1},...,c_{5}$ can be disregarded here, since they only change the angles of the rotational gates, but not the overall structure of the circuit.
Figure 8: Example circuit generated by the quantum walk state preparation algorithm presented in this work. The leading CX gates in the dashed rectangles can be removed without affecting the final state.
First, we start from the |$\ket{00001}$ state by applying an X gate to the last qubit. In the second segment of the circuit, we want to connect with the state |$\ket{00110}$. This state is different from the previous state in bits 2, 3, and 4. We arbitrarily choose qubit 2 as the target qubit and make the other bits consistent by applying CX gates to qubits 3 and 4, with control on qubit 2, which transforms the |$\ket{00110}$ state to |$\ket{00101}$. This enables the quantum walks between the states |$\ket{00001}$ and |$\ket{00101}$, which consist of the Rz gate that establishes the correct phase of $c_{1}$, and Rx gate, which establishes the correct magnitude of $c_{1}$. No controls are necessary on Rz and Rx gates here since only one state has non-zero amplitude at this point. After that, we apply the same CX gates again to transform the newly populated state back to |$\ket{00110}$.
In segment 3, we want to connect the states |$\ket{00110}$ and |$\ket{00111}$. These states are already adjacent, so no CX conjugation is necessary here. We want to make sure that Rz and Rx gates only affect |$\ket{00110}$ and not |$\ket{00001}$, so we add the minimally necessary number of controls to make it happen. In this case, it is sufficient to control on qubit 3. After these operations, the correct phase and magnitude of $c_{2}$ are established.
Continuing in the same fashion, we connect the states |$\ket{00111}$ and |$\ket{01001}$ in segment 4, and states |$\ket{01001}$ and |$\ket{01011}$ in segment 5. After this, all phases and magnitudes of the target coefficients are correct, except for the phase of $c_{5}$. To fix the phase of $c_{5}$, we choose the closest zero-amplitude state to the last populated state and interact with it via the Rz gate (no Rx gate is needed here). In this case, we chose the state |$\ket{11011}$ and control on qubits 2 and 3, where the control on qubit 2 is conjugated with X gates to implement a 0-control.
Note that the leading CX gates, marked with the dashed rectangles in Fig. 8 can be removed, since their control qubits, at the time of the gate application, are guaranteed to be in the 0-state. This optimization can only be done at the beginning of the circuit, before any CRx gates acted on the corresponding qubits, therefore we show these gates here for the purpose of demonstrating a general circuit structure.
V Performance
We tested the validity of our method by preparing 1000 random quantum states for each combination of $n$ and $m$ shown in Fig. 9. Each state was prepared using the quantum walk based techniques presented in this work with control reduction and SHP walk order. Additionally, the same states were also prepared with Qiskit’s built-in prepare_state method (which is based on Ref. [31]), shown here for comparison. All circuits have been transpiled into single-qubit gates and CX basis using Qiskit’s transpiler with optimization level 3.
As can be seen from Fig. 9, Qiskit’s performance scales exponentially and is very similar regardless of the value of $m$. In fact, for $m=n^{2}$ and $m=2^{n-1}$ Qiskit produces circuits with exactly the same number of CX gates regardless of the state being prepared.
In contrast to Qiskit, our method takes advantage of the sparsity of the target state and is expected to outperform Qiskit’s built-in method for any asymptotically sparse state (i.e. $m=\mathcal{O}(poly(n))$) for sufficiently large values of $n$. However, for asymptotically dense states (i.e. $m=\mathcal{O}(2^{n})$) our method is less efficient than Qiskit for sufficiently large values of $n$.
The exact code and data for these numerical experiments can be found in the linked repository (see Data Availability section Code and Data Availability).
Figure 9: The average number of CX gates necessary to prepare a random state using the uniformly controlled rotation method [31] used by Qiskit’s built-in state preparation algorithm and the quantum-walk based algorithm presented in this paper for different values of $n$ and $m$.
VI Conclusions and Future Directions
In this work, we developed an algorithm that can be used to convert CTQWs on dynamic graphs that consist only of self-loops or single edges to the quantum gate model. This algorithm serves as the basis for a deterministic state preparation routine that has complexity $\mathcal{O}(nm)$ where $n$ is the number of qubits and $m$ is the number of basis states in the target state. Furthermore, we introduce multiple methods that can reduce the CX count of the state preparation algorithm and test these methods against the uniformly controlled rotation method [31] used by the Qiskit transpiler. We find that our state preparation methods require fewer CX gates when the target state has a polynomial number of non-zero amplitudes.
The quantum walks QSP framework we present posts many interesting possible avenues for future investigation. As mentioned in Sections IV.2 and IV.3, the gate count for preparing an arbitrary state can be reduced depending on the order in which the state transfer occurs. Without control reduction, the optimal sequence of walks is given by the minimum spanning tree approach. Determining the optimal sequence of basis state transfers in the presence of control reduction is more complicated and could be an interesting direction for future research. Additionally, if a target state is comprised of basis states with symmetry, it may be possible to cleverly reduce the number of gates required to create the target state by exploiting symmetry. Another interesting avenue for future work is to compare against the methods in Refs. [6, 44].
Author contributions
AG developed the CTQWs to gates conversion, formulated control reduction, wrote code, gathered data, constructed proofs, and calculated complexities. RH developed the algorithm that converts CTQWs on dynamic graphs to the circuit model. CC wrote code, developed state preparation approaches, and tested circuit optimization techniques. IG developed the minimum spanning tree approach, CRz implementation of the self-loop walks, wrote code, collected data and prepared some figures. JL also developed the circuit optimization for reducing control bits. TT helped develop state preparation approaches and evaluation methodology. ZHS had the idea of creating a dictionary between single edge quantum walks and quantum circuits.
All authors wrote, read, edited, and approved the final manuscript.
Acknowledgements
The authors would like to thank Mostafa Atallah for useful conversations regarding this work.
AG, JL, and ZHS acknowledge DOE-145-SE-14055-CTQW-FY23. CC and TT acknowledge DE-AC02-06CH11357. IG and RH acknowledge DE-SC0024290. The funder played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
Competing interests
All authors declare no financial or non-financial competing interests.
Code and Data Availability
The code and data for this research can be found at https://github.com/GaidaiIgor/quantum_walks.
The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory (“Argonne”). Argonne, a U.S. Department of Energy Office of Science laboratory, is operated under Contract No. DE-AC02-06CH11357. The U.S. Government retains for itself, and others acting on its behalf, a paid-up nonexclusive, irrevocable worldwide license in said article to reproduce, prepare derivative works, distribute copies to the public, and perform publicly and display publicly, by or on behalf of the Government. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan. http://energy.gov/downloads/doe-public-access-plan.

### v2

@@
-thanks: Contributed equallythanks: Contributed equally
+Efficient Sparse State Preparation via Quantum Walks
 Abstract
 Continuous-time quantum walks (CTQWs) on dynamic graphs, referred to as dynamic CTQWs, are a recently introduced universal model of computation that offers a new paradigm in which to envision quantum algorithms.
-In this work we develop a mapping from dynamic CTQWs to the gate model of computation in the form of an algorithm to convert arbitrary single edge walks and single self loop walks, which are the fundamental building blocks of dynamic CTQWs, to their circuit model counterparts.
-We use this mapping to introduce an arbitrary quantum state preparation framework based on dynamic CTQWs.
+In this work we develop a mapping from dynamic CTQWs to the gate model of computation in the form of an algorithm to convert arbitrary single-edge walks and single self-loop walks, which are the fundamental building blocks of dynamic CTQWs, to their circuit model counterparts.
+We use this mapping to introduce an efficient sparse quantum state preparation framework based on dynamic CTQWs.
 Our approach utilizes global information about the target state, relates state preparation to finding the optimal path in a graph, and leads to optimizations in the reduction of controls that are not as obvious in other approaches.
 Interestingly, classical optimization problems such as the minimal hitting set, minimum spanning tree, and shortest Hamiltonian path problems arise in our framework.
-We test our methods against uniformly controlled rotations methods, used by Qiskit, and find ours requires fewer CX gates when the target state has a polynomial number of non-zero amplitudes.
+We test our methods against the state-of-the-art ancilla free sparse state preparation method and a uniformly controlled rotation method, which is used by Qiskit, and find ours requires fewer CX gates when the target state has a polynomial number of non-zero amplitudes.
 preprint: APS/123-QED
@@
 Furthermore, the authors of [20] found that CTQWs on at most three dynamic graphs can be used to implement the equivalent of a universal gate set.
-However, the opposite conversion from CTQWs to the gate model is less studied.
-Since CTQWs on dynamic graphs have yet to be implemented on hardware, converting them to the gate model is necessary to simulate them on existing hardware.
+While prior work has focused on representing quantum gate operations as CTQWs, the opposite conversion – generating a quantum circuit from a CTQW – is less well-studied.
+Since CTQWs on dynamic graphs have yet to be natively implemented on hardware, and given recent progress in the development of gate-based quantum hardware [21, 22, 23, 24], converting CTQWs to the gate model is necessary to execute them on existing quantum hardware.
 Additionally, developing an algorithm for such a conversion offers an alternative way of thinking about circuit design, which may result in simpler circuits, compiler optimization techniques, and new ansatzes for QAOA [25, 26, 27, 28], MA-QAOA [29, 30, 31] or VQE [32, 33, 34], where the trainable parameters correspond to dynamic graph propagation times.
-To this end, we develop an algorithm that converts CTQWs on single edge graphs and single self-loop graphs, which can serve as a basis for arbitrary graphs, to a sequence of gates in the circuit model.
+To this end, we develop an algorithm that converts CTQWs on single-edge graphs and single self-loop graphs, which can serve as a basis for arbitrary graphs, to a sequence of gates in the circuit model.
 This conversion algorithm is used as a foundation for a new deterministic arbitrary quantum state preparation (QSP) method, where the CX count of the resulting circuit is linear in the number of $m$ non-zero amplitude computational basis states that comprise the target state and the number of qubits $n$.
@@
 Our framework does not require ancillas.
-Figure 1: This figure lays out the framework for QSP via quantum walks.
-Alternating single edge quantum walks with self-loop quantum walks, allows arbitrary QSP.
-We provide conversion methods to construct the gate based circuit.
+Figure 1: Alternating single-edge and self-loop quantum walks yields a framework for arbitrary QSP.
+We provide conversion methods to construct the gate-based circuit.
 Control reduction is a method for reducing controls that we introduce in this paper.
@@
 As such, our method is best suited for the asymptotically sparse states ($m=\mathcal{O}(poly(n))$), but can, in principle, also work for the asymptotically dense states ($m=\mathcal{O}(2^{n})$).
-The walk framework that we present here is flexible, scalable, and intuitive, building upon well-established graph-based algorithms to efficiently prepare arbitrary quantum states.
+The walk framework that we present here is flexible, scalable, and intuitive, building upon well-established graph-based algorithms to efficiently prepare sparse quantum states.
 This paper is organized as follows.
@@
 II.
-We then introduce an algorithm that converts CTQWs on single edge dynamic graphs and single self loop graphs to the circuit model in Sec.
+We then introduce an algorithm that converts CTQWs on single-edge dynamic graphs and single self-loop graphs to the circuit model in Sec.
 III.
@@
 IV, we introduce a deterministic state preparation algorithm and analyze its circuit complexity.
-We compare our state preparation methods to the uniformly controlled rotation method [5, 31] used by Qiskit in Sec.
+We compare our state preparation methods to the sparse state preparation method from [6] and to the uniformly controlled rotation method [5, 35] used by Qiskit in Sec.
 V.
+In both cases, we found that our algorithm requires fewer CX gates.
 Finally, we discuss future work in Sec.
@@
 [39] provides an $\mathcal{O}(n^{2})$ depth divide-and-conquer method to prepare arbitrary states at the cost of needing $\mathcal{O}(2^{n})$ ancillary qubits.
+In Ref.
+[40], the ancilla overhead is improved to $\mathcal{O}(n)$.
 For arbitrary QSP, the state of the art deterministic protocol achieves $\mathcal{O}(\frac{23}{24}2^{n})$ CX scaling on even numbers of qubits [41].
@@
 This type of method is attractive for real-world applications of quantum computers because they can take advantage of sparsity while still remaining agnostic to particular details of the state.
-Another method using decision trees [43] has $\mathcal{O}(nm)$ CX gates, but it requires an ancilla qubit.
+Other methods achieve this, but require ancillas [48, 49].
+The method using decision trees [49] requires 1 ancilla qubit.
 The method in Ref.
@@
 III Conversion from CTQW on dynamic graphs to the Circuit Model
-A standard basis for any adjacency matrix is given by the set of adjacency matrices corresponding to all single edge walks and all single self-loop walks.
+A standard basis for any adjacency matrix is given by the set of adjacency matrices corresponding to all single-edge walks and all single self-loop walks.
 We introduce the conversion methods that can construct the gate-based representations for these walks on $n$ qubits.
 Throughout this paper, the considered states are represented in the standard computational basis set.
-III.1 Single Edge CTQW to Circuit Model
-The adjacency matrix for a single edge graph connecting basis states |$\ket{j}$ and |$\ket{k}$ is given by
+III.1 Single-Edge CTQW to Circuit Model
+The adjacency matrix for a single-edge graph connecting basis states |$\ket{j}$ and |$\ket{k}$ is given by
 $A(j,k)=\ket{j}\!\!\bra{k}+\ket{k}\!\!\bra{j}$ | (2)
@@
 Then apply the CRx($2t$) gate as before and apply the reverse of the conjugating CX sequence.
-An example conversion of a single edge walk to gates is shown in Fig.
+An example conversion of a single-edge walk to gates is shown in Fig.
 3.
@@
 The $(n-1)$-controlled CRx gate is a multi-controlled special unitary, and can be decomposed in $\mathcal{O}(n)$ CX and single-qubit gates [51].
-Figure 3: When a single edge walk is between states with non-unit Hamming distance from each other, CX gates might be required in addition to the CRx gate.
+Figure 3: When a single-edge walk is between states with non-unit Hamming distance from each other, CX gates might be required in addition to the CRx gate.
 III.2 Single Self-Loop CTQW to Circuit Model
@@
 [17] use walks on arbitrary graphs in their proof of universality.
-In this section we prove that using only single edge and single self-loop walks is sufficient to decompose an arbitrary unitary.
+In this section we prove that using only single-edge and single self-loop walks is sufficient to decompose an arbitrary unitary.
 Proposition 1.
-An arbitrary $d\times d$ unitary can be decomposed into a series of single self-loop and single edge CTQWs.
+An arbitrary $d\times d$ unitary can be decomposed into a series of single self-loop and single-edge CTQWs.
 Proof.
@@
 (3) and III.2 Eq.
-(6), a single edge CTQW corresponds to Rx’, and a sequence of two single self-loop CTQWs corresponds to W’ and Rz’.
-Thus, an arbitrary unitary can be decomposed into a series of single self-loop CTQWs and single edge CTQWs.
+(6), a single-edge CTQW corresponds to Rx’, and a sequence of two single self-loop CTQWs corresponds to W’ and Rz’.
+Thus, an arbitrary unitary can be decomposed into a series of single self-loop CTQWs and single-edge CTQWs.
 IV Deterministic state preparation via quantum walks
-The above conversions allow one to write single edge walk dynamics and single self-loops in terms of the circuit model.
+The above conversions allow one to write single-edge walk dynamics and single self-loops in terms of the circuit model.
 In this section, we describe how a series of such quantum walks can be used to prepare an arbitrary quantum state.
 IV.1 General Approach
-Given the task of preparing a quantum state with $m<2^{n}$ non-zero amplitudes on $n$ qubits, we start by presenting a high-level description of the algorithm to accomplish this task:
+Given the task of preparing a quantum state with $m<2^{n}$ non-zero amplitudes on $n$ qubits, we start by presenting a high-level overview of the algorithm to accomplish this task:
   1. 1.
-    Create a tree graph connecting the non-zero amplitudes, as shown in Fig. 4.
+    Construct a tree graph connecting the non-zero amplitudes of the target state in some way. See Fig. 4 for an example.
   2. 2.
-    Starting from an arbitrarily chosen root, and following a graph traversal order, perform a self-loop walk for each encountered node and a single edge quantum walk for each encountered edge.
+    Starting from some node, and following frontier-based graph traversal order (see Sec.
+IV.5), perform a self-loop walk for each encountered node and a single-edge quantum walk for each encountered edge.
   3. 3.
-    Convert the sequence of the quantum walks to the circuit representation.
-Initially, the system starts with all amplitude in the root state, which can be constructed from the ground |$\ket{0}$ state simply by applying the X gates to the necessary qubits.
-The graph traversal order can be be determined by any graph traversal algorithm, such as Depth First Search or Breadth First Search.
-An arbitrary amount of amplitude (absolute value) from the source can be transferred to the destination for any pair of the connected states corresponding to each single edge walk, and all non-zero amplitude states are connected transitively via a tree.
-Therefore, an arbitrary distribution of absolute values of amplitude (but not phases) can be established on the set of non-zero amplitude basis states via single edge walks.
+    Convert the sequence of the quantum walks to the circuit representation (which may include control reduction and CX forward- and backward-propagation, as described below).
+Frontier-based graph traversal order selects the next node to visit from a set of nodes adjacent to already visited nodes.
+Many different heuristics can be used to construct the sequence of walks in step 1 or choose a particular traversal order in step 2.
+Initially, the system starts with all amplitude in the root state, which can be constructed from the ground |$\ket{0}^{\otimes n}$ state by applying the X gates to the necessary qubits.
+An arbitrary amount of amplitude (absolute value) from the source can be transferred to the destination for any pair of the connected states corresponding to each single-edge walk, and all non-zero amplitude states are connected transitively via a tree.
+Therefore, an arbitrary distribution of absolute values of amplitude (but not phases) can be established on the set of non-zero amplitude basis states via single-edge walks.
 Each self-loop walk on a basis state |$\ket{z_{i}}$ establishes the correct phase of the corresponding coefficient $c_{i}$, but does not change the absolute value of it.
-Therefore, performing self-loop walks on each non-zero amplitude basis state establishes the correct phases of the corresponding coefficients, without interfering with the absolute values established by the single edge walk.
-Therefore, an arbitrary quantum state can be prepared with a sequence of single edge walks and self-loop walks, as described in the algorithm above.
-In the context of the state preparation, CRz gates can easily be used instead of CP gates to implement the self-loop walks, since every single edge walk transfers the amplitude to a new basis state with zero amplitude.
-Thus, as described in the previous section, the same basis state can be used to establish the correct phase on the source state for the single edge walk without additional CX conjugations.
+Therefore, performing self-loop walks on each non-zero amplitude basis state establishes the correct phases of the corresponding coefficients, without interfering with the absolute values established by the single-edge walk.
+Therefore, an arbitrary quantum state can be prepared with a sequence of single-edge walks and self-loop walks, as described in the algorithm above.
+In the context of the state preparation, CRz gates can easily be used instead of CP gates to implement the self-loop walks, since every single-edge walk transfers the amplitude to a new basis state with zero amplitude.
+Thus, as described in the previous section, the same basis state can be used to establish the correct phase on the source state for the single-edge walk without additional CX conjugations.
 However, this will not apply to the leaf nodes (degree one vertices) in the tree, which will require another zero-amplitude state to interact with.
@@
 Let $n$ denote the number of qubits and $S=\{\ket{z_{1}},\ket{z_{2}},...,\ket{z_{k}}\}$ denote the set of visited nodes.
-Suppose we wish to perform a single edge walk from |$\ket{z_{j}}\in S$ to |$\ket{z_{\ell}}\notin S$ and |$\ket{z_{j}}$ and |$\ket{z_{\ell}}$ have a Hamming distance of 1.
+Suppose we wish to perform a single-edge walk from |$\ket{z_{j}}\in S$ to |$\ket{z_{\ell}}\notin S$ and |$\ket{z_{j}}$ and |$\ket{z_{\ell}}$ have a Hamming distance of 1.
 Let $b$ denote the qubit where |$\ket{z_{j}}$ is different from |$\ket{z_{\ell}}$, i.e.
 $z_{j}[b]\neq z_{\ell}[b].$ Let $D=\{\{k\,|\,z_{i}[k]\neq z_{j}[k],k\neq b\}\,|\,z_{i}\in S\setminus z_{j}\}$ denote the set of differing bits of the visited nodes with $z_{j}$ excluding the qubit $b$.
-For the gate-based representation of a single edge walk from basis state |$\ket{z_{j}}$ to |$\ket{z_{\ell}}$, it is sufficient to control the CRz or CRx gate on any hitting set of $D$, where the values of controls are equal to the corresponding bits of $z_{j}$.
+For the gate-based representation of a single-edge walk from basis state |$\ket{z_{j}}$ to |$\ket{z_{\ell}}$, it is sufficient to control the CRz or CRx gate on any hitting set of $D$, where the values of controls are equal to the corresponding bits of $z_{j}$.
 A hitting set of $D$ is a collection of elements $h$ such that $h\cap d_{i}\neq\emptyset$ for all $d_{i}\in D$.
@@
 Then we apply the previously described control reduction method to $\tilde{S}$.
+Note that in this case, any of the differing qubits can be used as a target for the CRz or CRx gates, with the appropriate CX conjugation.
+However, the target qubit cannot be chosen as a control, therefore the choice of target affects the efficiency of control reduction.
+In the following sections, when a target qubit is not explicitly dictated by a particular walk order, we consider all options for the target and choose the one that results in the minimum number of controls.
 In order to minimize the number of controls, one needs to minimize the size of the hitting set, i.e.
-solve the minimum hitting set problem, which is known to be NP-complete.
+solve the minimum hitting set (MHS) problem, which is known to be NP-complete.
 As such, no polynomial algorithm is known to solve it exactly, but a number of heuristics exist that can provide good suboptimal solutions quickly [54, 55].
@@
 5).
+The MHS problems that we consider in the following sections are solved approximately.
+Example: Consider the sequence of walks [[001, 111], [111, 110]].
+Suppose we have implemented the first walk and are now implementing the second walk.
+The second walk transfers amplitude from 111 to 110.
+The current state is a superpostion of |$\ket{001}$ and |$\ket{111}$.
+Thus, we want the walk to affect only the amplitude of basis state 111 and leave 001 alone.
+Therefore, we need the controls for the Rx gate to be satisfied only by 111.
+Controlling the CRx gate on qubit 0 accomplishes this.
 Figure 5: The average number of CX gates necessary to implement a random state with $n$ qubits with and without control reduction.
 Each state consist of $m=n$ non-zero amplitudes and each point is the average over 1000 random states.
-IV.3 Walk Order
-Any tree constructed in the first step is, in principle, sufficient for state preparation but not all trees result in the same CX gate count.
-As mentioned in the previous section, the cost of a walk between a given pair of basis states depends on the Hamming distance between them.
-Therefore, connecting the basis states that are close to one another (in terms of their Hamming distance) minimizes the number of CX gates in the resulting circuit.
-Let us consider some of the more advantageous options.
+IV.3 CX Forward- and Backward-Propagation
+As described in Sec.
+III.1, single-edge quantum walks between states with Hamming distance greater than one require conjugation with CX gates on both sides of the CRx gate.
+The CX gates on the left side bring the desired basis states to a Hamming distance of 1, where they interact via the CRx gate, and the ones on the right move the basis states back to their original positions.
+Rather than immediately moving the states back after the application of the CRx gate, one can instead keep working with the transformed basis and move the right branch of the conjugating CX gates to the end of the circuit, which allows many of those CX gates from different single-edge walks to cancel.
+This is called forward-propagation of CX gates.
+The backward-propagation of CX gates is the same method applied to the reverse of the circuit.
+The advantage of the backward-propagation is that the CX gates are moved to the beginning of the circuit.
+Since the initial state is the ground state, these CX gates act trivially and can be removed.
+Example: Consider the sequence of walks [[001, 111], [011, 110]].
+The first walk [001, 111] has a Hamming distance of 2.
+For target qubit 1, we require CX10 to bring the states to a Hamming distance of 1.
+Applying this to the basis states we get [001, 011], [111, 010].
+After applying the multicontrol gate(s) for the first walk, we do not rotate the basis states back to the original frame, but instead keep track of the CX rotation which will be applied at the end of the circuit.
+This is an example of CX forward-propagation.
+IV.4 Leaf Node Considerations
+When a walk is performed between two non-leaf nodes of the tree (see Fig.
+4), we use a CRz gate to adjust the phase on the walk origin node, followed by a CRx gate to transfer the amplitude to the destination node, both of which have the same control set, since CRz operation does not change the number of non-zero amplitude basis states.
+The phase on the destination node is adjusted once we consider the first walk out of that node.
+However, when the destination node is a terminal (leaf) node, there are no other walks that start from it and we still need to fix its phase.
+One way to solve this issue is to use a separate CRz gate, but it adds another multi-controlled gate to the circuit and the number of controls for it is likely to be greater compared to the walk into the leaf node due to the fact that we will need to distinguish the leaf from a larger number of non-zero amplitude states and control reduction will be less efficient.
+An alternative approach is to combine this last CRz gate with the CRz and CRx gates from the walk leading into the leaf, as one gate that simultaneously transfers the probability amplitude and adjusts the phase on both ends of the walk.
+This reduces the number of multi-controlled operations from 3 to 1, but the downside of this approach is that the resulting gate will not belong to SU(2) group.
+As such, the number of CX gates in its decomposition will be proportional to $\mathcal{O}(n^{2})$, instead of $\mathcal{O}(n)$ necessary for the regular SU(2) gates, such as CRz or CRx.
+[56] Asymptotically, it makes this approach less efficient for sufficiently large values of $n$, but for our dataset it is still favorable.
+IV.5 Tree Construction and Walk Order
+Any tree constructed in the first step is, in principle, sufficient for state preparation, but not all trees result in the same CX gate count.
+Moreover, different walk orders on the same tree can can result in different number of CX gates as well.
+In this section we consider some of the options for tree construction and walk order.
+Best Orders. We first describe heuristics that typically outperform state-of-the-art methods on average.
+MHS Nonlinear.
+As demonstrated in section Sec.
+IV.2, control reduction is a powerful technique that allows to significantly simplify the state preparation circuit.
+The efficiency of it depends on how easily a particular state can be separated from the rest of the currently existing states with non-zero amplitudes, which in turn depends on the local density of states around the state that we want to separate.
+The basic idea of the following walk order heuristic is that to maximize the efficiency of control reduction, we need to walk through hard-to-separate states first, before the density of states around them increases, which allows us to separate them using smaller number of controls.
+Since the target qubit has a significant effect on the CX count, it is determined during the walk order construction as well.
+More specifically, we construct a sequence of walks in reverse time order, from last to first.
+Given a list of basis states, we choose the state that requires the smallest number of controls to differentiate it from the rest of the basis states by solving the Minimum Hitting Set (MHS) problem.
+When two states are equally easy to distinguish, we choose the state with the largest number of bits that are different from the rest.
+Let us call this state $z_{1}$.
+The target qubit $t$ is the MHS bit with the smallest frequency.
+The set differentiated by $t$ is then ordered by the easiest element to distinguish via the MHS.
+When two elements are equally easy to distinguish, we order them by Hamming distance from $z_{1}$.
+The easiest element is set as $z_{2}$.
+The ($z_{1}$, $z_{2}$, $t$) that requires the fewest controls is added to the path.
+The number of controls required is upper bounded by the size of the MHS for $z_{1}$ plus the size of the MHS for $z_{2}$.
+All the elements are then updated by the back-propagation of CX gates.
+$z_{2}$ is removed from the list and the process is repeated until every node is visited by the path.
+The circuit is constructed from the path of CTQWs in terms of the original basis states and target qubits and with CX backward-propagation (see Alg.
+2 for pseudocode).
+MHS Linear.
+This protocol follows the same procedure as MHS Nonlinear except that each step we assign $z_{1}$ of the previous pair as the $z_{2}$ for the next pair.
+This results in a linear graph.
+In contrast, in MHS Noninear $z_{2}$ of the next pair is chosen using the MHS methods.
+Greedy Insertion.
+The linear greedy algorithm takes an ordered list of basis states and sequentially finds the best position to insert the next basis state.
+The classical complexity of this algorithm is $\mathcal{O}(nm^{4}+C)$, where $C$ is the classical complexity of the method that generates the initial ordering.
+Typically, the greedy insertion performs better when given an initial ordering that already performs well (see Alg.
+3 for pseudocode).
+Other Orders. The following are other orders we considered that on average typically do not outperform state-of-the-art methods.
 Minimum Spanning Tree.
@@
 Shortest Hamiltonian Path.
-MST approach minimizes the total Hamming distance for the walks, which corresponds to the optimal walk order if control reduction is not taken into account.
+The MST approach minimizes the total Hamming distance for the walks, which corresponds to the optimal walk order if control reduction is not taken into account.
 However, when control reduction is applied, other walk orders may be more efficient since they may allow for additional control reduction.
-In general, it is difficult to calculate in advance which walk orders are better for the maximum control reduction, but from the numerical experiments we discovered that, on average, path graphs (trees with exactly two leaves) are more amenable to it than other trees.
 Therefore, instead of MST, one could find the Shortest Hamiltonian Path (SHP) in the same graph of the target basis states as used to build the MST.
@@
 For example, for a fully dense state, the average Hamming distance given by the sorted walk order is 2, which is much better than $n/2$ achieved by a random walk order.
-A comparison of the different walking orders presented in this section for the case of $m=n$ with control reduction is shown in Fig.
+A comparison of the different walking orders presented in this section for the case of $m=n$ is shown in Fig.
 7.
-We empirically validate the four quantum walk order methods and find that on the prepared states, SHP requires the fewest CX gates, although it approaches the MST method as $n$ increases.
+We empirically validate the quantum walk order methods and find that on the prepared states, the MHS methods and greedy methods require the fewest CX gates.
+Greedy(MHS) Combined is the greedy method on the initial ordering of MHS Linear and the better circuit between the two methods is used.
 (a) MST for a sparse state.
@@
 6d and  6b, the costs of the edges are suppressed except for the path.
-Figure 7: The average number of CX gates necessary to implement a random state with given number of qubits for the considered walk orders.
-Each state consist of $m=n$ non-zero amplitudes and each point is the average over 1000 random states.
-IV.4 Additional Optimizations
+IV.6 Additional Optimizations
 A single edge between states |$\ket{j}$ and |$\ket{k}$ corresponds to a CRx gate, which introduces an imaginary phase on |$\ket{k}$.
@@
 We plan to implement this in our code in the future.
-IV.5 Example
-Suppose we want to prepare the state
-[…]
-Using the algorithm described above, with sorted walk order and control reduction, one can generate a circuit shown in Fig.
-8.
-The exact values of the coefficients $c_{1},...,c_{5}$ can be disregarded here, since they only change the angles of the rotational gates, but not the overall structure of the circuit.
-Figure 8: Example circuit generated by the quantum walk state preparation algorithm presented in this work.
-The leading CX gates in the dashed rectangles can be removed without affecting the final state.
-First, we start from the |$\ket{00001}$ state by applying an X gate to the last qubit.
-In the second segment of the circuit, we want to connect with the state |$\ket{00110}$.
-This state is different from the previous state in bits 2, 3, and 4.
-We arbitrarily choose qubit 2 as the target qubit and make the other bits consistent by applying CX gates to qubits 3 and 4, with control on qubit 2, which transforms the |$\ket{00110}$ state to |$\ket{00101}$.
-This enables the quantum walks between the states |$\ket{00001}$ and |$\ket{00101}$, which consist of the Rz gate that establishes the correct phase of $c_{1}$, and Rx gate, which establishes the correct magnitude of $c_{1}$.
-No controls are necessary on Rz and Rx gates here since only one state has non-zero amplitude at this point.
-After that, we apply the same CX gates again to transform the newly populated state back to |$\ket{00110}$.
-In segment 3, we want to connect the states |$\ket{00110}$ and |$\ket{00111}$.
-These states are already adjacent, so no CX conjugation is necessary here.
-We want to make sure that Rz and Rx gates only affect |$\ket{00110}$ and not |$\ket{00001}$, so we add the minimally necessary number of controls to make it happen.
-In this case, it is sufficient to control on qubit 3.
-After these operations, the correct phase and magnitude of $c_{2}$ are established.
-Continuing in the same fashion, we connect the states |$\ket{00111}$ and |$\ket{01001}$ in segment 4, and states |$\ket{01001}$ and |$\ket{01011}$ in segment 5.
-After this, all phases and magnitudes of the target coefficients are correct, except for the phase of $c_{5}$.
-To fix the phase of $c_{5}$, we choose the closest zero-amplitude state to the last populated state and interact with it via the Rz gate (no Rx gate is needed here).
-In this case, we chose the state |$\ket{11011}$ and control on qubits 2 and 3, where the control on qubit 2 is conjugated with X gates to implement a 0-control.
-Note that the leading CX gates, marked with the dashed rectangles in Fig.
-8 can be removed, since their control qubits, at the time of the gate application, are guaranteed to be in the 0-state.
-This optimization can only be done at the beginning of the circuit, before any CRx gates acted on the corresponding qubits, therefore we show these gates here for the purpose of demonstrating a general circuit structure.
 V Performance
-We tested the validity of our method by preparing 1000 random quantum states for each combination of $n$ and $m$ shown in Fig.
-9.
-Each state was prepared using the quantum walk based techniques presented in this work with control reduction and SHP walk order.
-Additionally, the same states were also prepared with Qiskit’s built-in prepare_state method (which is based on Ref.
-[31]), shown here for comparison.
-All circuits have been transpiled into single-qubit gates and CX basis using Qiskit’s transpiler with optimization level 3.
-As can be seen from Fig.
-9, Qiskit’s performance scales exponentially and is very similar regardless of the value of $m$.
+In this section we present the comparison between the results of numerical simulations obtained for different walk orders of our single-edge state preparation method, the Merging States (MS) method introduced in [6] and Qiskit’s built-in state preparation method based on uniformly controlled rotations [5].
+The error bars in the figures denote the 95% confidence interval for the estimate of the mean.
+V.1 Comparison of Walk Orders
+First, we compare the performance of different walk orders presented in Sec.
+IV.5, which is shown in Fig.
+7.
+The greedy methods use the initial ordering specified in parenthesis.
+For the “Greedy Combined” results, we consider both Greedy and the method in parenthesis and choose the one with the lower CX count.
+As seen from the figure, all considered walking orders have roughly similar performance, with Random, Sorted, SHP and MST being somewhat worse than others.
+MHS Linear and Nonlinear outperform Greedy(Sorted) on average except at 5 qubits where only MHS Linear beats greedy.
+As expected, Greedy(MHS) Combined performs the best on average.
+Figure 7: Comparison of walk orders with $m=n$. Greedy(MHS) (Greedy(Sorted)) uses MHS Linear (Sorted) as the initial ordering. Each data point is averaged over 1000 random states.
+(a) $m=n$
+(b) $m=n^{2}$.
+Figure 8: CX count comparison between single-edge method with the best walk order (MHS Nonlinear) presented in this paper, state-of-the-art sparse state preparation method (Merging States) and Qiskit’s built-in state preparation method for (a) $m=n$ and (b) $m=n^{2}$.
+Each data point is averaged over 1000 randomly generated states.
+V.2 Comparison to Merging States and Qiskit
+In contrast to our single-edge state preparation method, which starts from one state and gradually splits existing states to form new states, the MS method of Ref.
+[6] instead focuses on the inverse process of merging basis states.
+The MS method can be interpreted as a series of quantum walks with specified targets for the multi-control gates and optimized with the CX backward-propagation method.
+Note that in MS, the authors of Ref.
+[6] use a single multi-controlled gate, denoted CG, to implement the merging (splitting) and the phase correction in one step.
+However, CG can also be written as a product of CP and CRx gates.
+In Figs.
+8a and 8b, we compare the performance of our single-edge state preparation method with the best walk order (MHS Nonlinear) to that of the Merging States method and Qiskit’s built-in prepare_state method (which is based on Ref.
+[35]).
+Just as in the previous section, an average CX gate count for all methods is calculated for a set of 1000 randomly generated sparse states ($m=n$ and $m=n^{2}$ non-zero amplitude basis states) for every value of $n$ from 5 up to 11.
+As can be seen from Figs.
+8a and 8b, Qiskit’s performance scales exponentially and is very similar regardless of the value of $m$.
 In fact, for $m=n^{2}$ and $m=2^{n-1}$ Qiskit produces circuits with exactly the same number of CX gates regardless of the state being prepared.
@@
 $m=\mathcal{O}(poly(n))$) for sufficiently large values of $n$.
-However, for asymptotically dense states (i.e.
-$m=\mathcal{O}(2^{n})$) our method is less efficient than Qiskit for sufficiently large values of $n$.
+Our methods outperforms not only sparsity-agnostic Qiskit’s method, but also the state-of-the-art sparse state preparation method – Merging States.
+As one can see from the plots, for all values of $n$, our method requires fewer CX gates.
 The exact code and data for these numerical experiments can be found in the linked repository (see Data Availability section Code and Data Availability).
-Figure 9: The average number of CX gates necessary to prepare a random state using the uniformly controlled rotation method [31] used by Qiskit’s built-in state preparation algorithm and the quantum-walk based algorithm presented in this paper for different values of $n$ and $m$.
-VI Conclusions and Future Directions
+VI Conclusions and Outlook
 In this work, we developed an algorithm that can be used to convert CTQWs on dynamic graphs that consist only of self-loops or single edges to the quantum gate model.
 This algorithm serves as the basis for a deterministic state preparation routine that has complexity $\mathcal{O}(nm)$ where $n$ is the number of qubits and $m$ is the number of basis states in the target state.
-Furthermore, we introduce multiple methods that can reduce the CX count of the state preparation algorithm and test these methods against the uniformly controlled rotation method [31] used by the Qiskit transpiler.
+Furthermore, we introduce multiple methods that can reduce the CX count of the state preparation algorithm and test these methods against the MS method of [6] and uniformly controlled rotation method [35] used by the Qiskit transpiler.
 We find that our state preparation methods require fewer CX gates when the target state has a polynomial number of non-zero amplitudes.
-The quantum walks QSP framework we present posts many interesting possible avenues for future investigation.
-As mentioned in Sections IV.2 and IV.3, the gate count for preparing an arbitrary state can be reduced depending on the order in which the state transfer occurs.
+In Ref.
+[49], the authors present another sparse state preparation method that requires only 1 ancilla.
+We did not directly compare to their method, but their protocol is outperformed by the MS method for low-degree polynomials $m(n)$ [49].
+Since our method outperforms the MS method in these regimes, our method should outperform theirs as well.
+The single-edge framework we present here posts many interesting possible avenues for future investigation.
+As mentioned in Sections IV.2 and IV.5, the gate count for preparing an arbitrary state can be reduced depending on the order in which the state transfer occurs.
 Without control reduction, the optimal sequence of walks is given by the minimum spanning tree approach.
@@
 Additionally, if a target state is comprised of basis states with symmetry, it may be possible to cleverly reduce the number of gates required to create the target state by exploiting symmetry.
-Another interesting avenue for future work is to compare against the methods in Refs.
-[6, 44].
 Author contributions
@@
 TT helped develop state preparation approaches and evaluation methodology.
-ZHS had the idea of creating a dictionary between single edge quantum walks and quantum circuits.
+ZHS had the idea of creating a dictionary between single-edge quantum walks and quantum circuits.
 All authors wrote, read, edited, and approved the final manuscript.
-Acknowledgements
+Acknowledgments
 The authors would like to thank Mostafa Atallah for useful conversations regarding this work.
@@
 The code and data for this research can be found at https://github.com/GaidaiIgor/quantum_walks.
+1:Convert from CTQWs to a circuit.
+2:Input:
+3:$walks$: The set of walks.
+4:$state$: The target state.
+5:Output:
+6:The circuit that implements the quantum walks.
+7:
+8:$walks\leftarrow$ Reverse of the walks.
+9:$circ\leftarrow$ Initial empty quantum circuit.
+10:for $z_{1}$, $z_{2}$, $target$, $pangle$, $xangle$ in walks do
+11:   if $target$ is not specified then
+12:      $target\leftarrow$The qubit index that results in the smallest number of controls.}
+13:   end if
+114:   {Apply the CX gates required to bring $z_{1}$ and $z_{2}$ to a Hamming distance of 1.}
+15:   Determine $mhs$ for $z_{1},z_{2},target$.
+16:   if $z_{2}$ is a leaf node then
+17:      {Apply controlled U Gate to $circ$ (corresponds to the phase, Rx, phase gate pattern described in Section IV.4.)}
+18:   else
+19:      Apply controlled SU(2) Gate to $circ$.
+20:   end if
+21:   Update the walks with the CX gates.
+22:end for
+23:return inverse of $circ$.
+Algorithm 1 Conversion
+1:Input:
+2:$basis$: A set of basis states to walk through.
+3:Output:
+4:A sequence of single-edge quantum walks [[$z_{1}$, $z_{2}$, $target$]i] through the provided basis states that minimizes the number of controls needed to implement the corresponding multi-controlled operations.
+5:
+6:$path\leftarrow$ empty.
+7:$mutable\_basis\leftarrow$ copy $basis$.
+18:while size of $basis>1$ do
+9:   $all\_mhs\_z_{1}\leftarrow$MHS for every element in $mutable\_basis$.}
+10:   $all\_diffs\leftarrow$for every element $z_{i}$ in $mutable\_basis$ get the subset $k_{i}=mutable\_basis/z_{i}$ and construct a list of list of the bits that differ with $z_{i}$.}
+11:   $all\_targets\leftarrow$For every element $mhs$ in $all\_mhs\_z_{1}$ get the smallest frequency index in the corresponding $all\_diffs[mhs]$}.
+12:   $all\_basis\_z_{2}\leftarrow$A list of list.
+For every element $z_{i}$ in $mutable\_basis$ get the subset $k_{i}\subset mutable\_basis/z_{i}$ such that $k_{i}$ is only covered by the hitting set $all\_mhs\_z_{1}[z_{i}]$ with the corresponding target $all\_targets[z_{i}]$.}
+13:   $all\_mhs\_z_{2}\leftarrow$A list of list of list. For every element $basis_{i}$ in $all\_basis\_z_{2}$ get the MHS for each element in $basis_{i}$.}
+14:   {Append to $path$ $[z_{1},z_{2},target]$ that corresponds to the smallest size($all\_mhs\_z_{1}[z_{1}]$)$+$size($all\_mhs\_z_{2}[z_{1}][z_{2}]$).
+Note that $z_{1}$ and $z_{2}$ are in terms of the original basis.}
+15:   Remove $z_{2}$ from $basis$.
+16:   Remove $z_{2}$ from $mutable\_basis$.
+217:   Update $mutable\_basis$ with CX gates.
+18:end while
+19:return the inverse of $path$.
+Algorithm 2 MHS Nonlinear
+1:The walks considered here all form linear graphs. Greedy finds the best position to insert the next basis state into the path sequence.
+2:Input:
+3:$basis$: Basis in some order.
+4:Output:
+5:The best circuit found.
+16:
+7:$path\leftarrow$ Empty list
+8:for $z$ in $basis$ do
+9:   $cxcount\leftarrow$ None.
+10:   $curr\_path\leftarrow$ None.
+11:   for $idx$ in (length($path$)+1) do
+12:      $temp\_path\leftarrow$ copy $path$.
+13:      Insert $z$ in $temp\_path$ at $idx$.
+14:      $temp\_cxcount\leftarrow$ CX count for constructed circuit with $temp\_path$.
+15:      if $cxcount$ is None or $temp\_cxcount<cxcount$ then
+16:         $curr\_path\leftarrow temp\_path$.
+17:      end if
+18:   end for
+19:   $path\leftarrow curr\_path$.
+20:end for
+21:return circuit for $path$.
+Algorithm 3 Greedy Insertion
 The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory (“Argonne”).